The answer to this is 1 minus the probability that they will have 3 or fewer children. This would happen only if they had a boy as the first, second or third child. The probability they have a boy as first child is 0.5
The probability they have a boy as second is 0.25
The probability they have a boy as third is 0.125
Thus the total probability is 0.875
And so the probability they will have more than three children is 1-0.875 or 0.125
What are the odds of being mauled by a tiger?
It likely depends on the tiger, and perhaps how hungry it is.
What is the probability of rolling an 8 with one roll of a single number cube?
If the cube has a 1,2,3,4,5,6,7,8, and a 9 on it, then you have 1/10 chance, or 10% because there is one that you want out of ten possibilities. If it has different digits, it would be 1/howmany digits there are, assuming that there is only one face with an eight.
How is probability used in biology?
It's used commonly to estimate the traits of a child of two parents. For example, the probability of the child having blue eyes, or curly hair, or even having genetic disease.
The probability that two cards drawn from a deck of cards are both aces?
The probability of drawing two Aces from a standard deck of 52 cards is 4 in 52 times 3 in 51, or 12 in 2652, or 1 in 221, or about 0.00452.
Define the term factorial in probability?
Often, when evaluating theoretical probability of something, say a lottery game or a card game, you consider permutations and combinations, in order to assess the number of outcomes that are possible versus the number that are desired.
When you work with permutations and combinations, you often work with factorials. In short, the factorial of a number is the product of that number and all of the numbers successively smaller by one until you get to one, i.e. 5 factorial is 5 x 4 x 3 x 2 x 1 or 120.
What is the probability of rolling a number less than 3 five times in a row on a number cube?
Normal die has 1,2,3,4,5,6
Less than 3 is thus 1,2 = i.e. 2 possibilities out of 6. 2/6 = 1/3
five times is thus 1/3 x 1/3 x 1/3 x 1/3 x 1/3 = 1/243
So probability is 1/243
Can you put in math questions on here and get an answer?
Yes, but it would help if you properly categorize the question. Also, know that most people resist answering what looks like homework, although some will answer it and try to provide an explanation so you can still learn the lesson. It depends on context.
For any particular card, (eg an 8) there are 4 such cards in a pack of 52. So the probability of drawing one of these is 4/52, or 1 in 13. After restoring the card to the pack and doing it again the probability is a further 1/13. So the overall probability of doing the 2 things in succession is 1/13 x 1/13 = 1/169.
How does experimental probability 43 out of 53 compare with theoretical probability 0.50?
43/53 = 0.81 : This is not even close to the theoretical probability of 0.50 ; With 0.50, the expected outcome is 26.5 out of 53. I would expect not to get 26 or 27, though - but a range from 21 to 33 (out of 53) would be in my expectations.
Imagine coin tosses (probability = 0.5): it is possible to get 43 out of 53, but the chance of this happening is very very small. I don't remember the exact formulas, but consider this: flip a coin 5 times. There are 32 possible outcomes. Only 6 of these 32 would give a 1 out of 5 (or less). This is 6/32 or about 19%. So to get 40 (or more) out of 50, you would need to get 4 (or more) out of 5, ten times, or (6/32)^10 = about 1 in 18 million. To get 43 (or more) out of 53 would be even a smaller chance than this. This is on the scale of chances of winning the lotto (definitely possible, but not probable).
How do you find binomial probabilities using a binomial table?
Read the instructions that accompany the table: they do not all have exactly the same layout.
Picking the first king: 4/52
Picking the second king: 3/51
Picking the queen of hearts: 1/50
Probability of the whole 3-step maneuver: (4/52) x (3/51) x (1/50) = 12 / 132,600 = 0.00905% (rounded)
What is the probability of rolling a five with a pair of dice?
There are 36 possible outcomes when we roll a pair of dice. Rolling a five can be done the following ways: 3-2, 4-1, 2-3 and 1-4. That's four ways to roll a five with one roll of the dice. The odds of rolling a five are 4 in 36, or 1 in 9. That makes the probability of rolling a five equal to 0.11111....
We know that the probability of something is a ratio of the number of times it can be done to the total number of things that are possible. Probabilities range from zero to one, inclusive. If something has a probability of zero, it cannot happen. Like rolling a 13. A 13 cannot be rolled with one roll of a pair of standard dice. Thus the probability of rolling a 13 is zero, as it cannot happen.
If something has a probability of one, it must happen. Like the probability of rolling an odd or an even number with one roll of a pair of dice. Since any number we roll must be either odd or even, then the probability of rolling an odd or an even number with one roll of a pair of dice is one. An odd or an even number must turn up, and that is why the probability of rolling an odd or an even number with one roll of the dice is one. Lastly, we've seen that the probability of rolling a given number with one roll of the dice is the number of times that number might appear divided by the total number of possibilities that might appear.
Is it possible for 2 homozygous recessive individuals to have a homozygous dominant child?
No, homozygous recessive indivuals MUST have two of the reccessive gene (bb), the possible combinations are bb,bb,bb,bb thus only a recessive trait is possible
What is the probability of drawing 4 aces in four draws?
4/52*3/51*2/50*1/49=24/6,497,400
It can be reduced.
1/270,725
Notice the numerator is the number of the desirable choices.
The denominator is the total number available.
What is the probability that the first die will have the same number as the second die?
On six sided dice: 1 out of 6. If you make a chart:
- 1 - 2 - 3 - 4 - 5 - 6
1 11 12 13 14 15 16
2 21 22 23 24 25 26
3 31 32 33 34 35 36
4 41 42 43 44 45 46
5 51 52 53 54 55 56
6 61 62 63 64 65 66
See that 6 out of the 36 possibilities (or 1 out of 6) have doubles, meaning that the second die is the same as the first die.
Why can't you use FOIL to multiply a binomial?
you can. i am in algebra II and that's what i was taught
Three coins are tossed simultaneously the probability of getting atmost one head?
1/2 or 50%
(1/8 prob of 0H + 3/8 prob of 1H)
What is an advantage of stereotactic radiosurgery?
The beam can be focused on any structure in the brain, allowing access to tumors or malformed blood vessels that cannot be reached by open-skull surgery.
What is the number of elements in an event space for rolling an even number on a single die?
The even numbers that can be rolled on a single die are 2,4, and 6 so the number of elements in the event space is 3. X={2,4,6}.
What is the probability of flipping 4 fair coins 1 time and getting all tails?
The probability of getting all tails when you flip 4 fair coins 1 time is:
P(4 tails) = (1/2)4 = 1/16 = 0.0625 = 6.25%
If you were throwing one numbered cube (dice) with each face numbered 1 to 6, then the chance of getting a number greater than two is 4/6 (or 2/3) since there are 4 numbers greater than 2, and six numbers in total.
Similarly, if you were throwing one dice, the chance of getting a number less than 2 is 1/6.
To determine the probability of getting both of these outcomes on two dice (greater than 2 on one dice and less than 2 on the other) we simply multiply these two values together.
2/3 x 1/6 = 2/18 = 1/9.
So there is a one in nine chance of this outcome occuring.